OpenStudy (countrygirl1431):

How many subsets does set A have if the set A has 3 elements? 3 6* 8 is this the correct answer

OpenStudy (zzr0ck3r):

If the set \(A\) is finite, then it has \(2^{|A|}\) in its power set.

OpenStudy (zzr0ck3r):

Maybe you are forgetting about the entire set, and the empty set?

OpenStudy (lynfran):

{a,b,c} the is subset are {a}, {b}, {c}, {a,b}, {a,c}, {b,c}, {a,b,c}, {}

OpenStudy (anonymous):

if n is the no. of elements then no. of subsets=\[2^n\]

OpenStudy (countrygirl1431):

im confused

OpenStudy (zzr0ck3r):

if a set has \(10 \) elements then its power set has \(2^{10}\) elements

OpenStudy (anonymous):

no. of elements in the set=3 no. of subsets=\[2^3\]

OpenStudy (zzr0ck3r):

if a set has \(4\) elements then its power set has \(2^4\) elements. If a set has \(3\) elements, then its power set has \(2^?\) elements

OpenStudy (countrygirl1431):

oh so it would be 3

OpenStudy (zzr0ck3r):

read what we wrote.

OpenStudy (anonymous):

\[2^3=2*2*2=?\]